Kiwidale Dairy is considering purchasing a new ice-cream maker. Two models, Smoothie and Creamy, are available and their information is given below (all costs and profits are in dollars):

Smoothie Creamy
First Cost 13,500 35,000
Service Life 14 Years 12 Years
Annual Profit 4,350 11,050
Annual Operating Cost 1,200 3,475
Salvage Value 2,000 5,100

Which alternative is better if the MARR for Kiwidale Dairy is 18.3%. Assume that each alternative can be repeated indefinitely {Perform all calculations using 5 significant figures and round any monetary answers to the nearest cent}.

a.) What is the Annual Worth of the Smoothie Model?

b.) What is the Annual Worth of the Creamy Model?

c.) Which model would you choose?
Answer #1

EUAW of smoothie model = -13500 * (A/P, 18.3%,14) + 4350 - 1200 + 2000 * (A/F, 18.3%,14)

= -13500 * 0.202234 + 4350 - 1200 + 2000 * 0.019234

= 458.31

EUAW of creamy model = -35000 * (A/P, 18.3%,12) + 11050 - 3475 + 5100 * (A/F, 18.3%,12)

= -35000 * 0.211097 + 11050 - 3475 + 5100 * 0.028097

= 329.89

As the EUAW of smoothie model is more it should be selected